# Multi-group Uncertainty Quantification for Long-form Text Generation

### v1

@@ 4.2 Collecting group features
  • nationality: Following Min et al. (2023a), who use nationality derived from Wikidata to sample their dataset of human entities, we take the property country of citizenship (or place of birth when not available) and categorize the corresponding value into the following categories defined by Min et al. (2023a): Asia/Pacific, Europe/Middle East, North America, Latin/South America/Africa.
@@
To form groups $\mathcal{G}$ from these attributes, we take all $1$ and $2$-way combinations [^5] To reduce computation, we train the quantile regression models in CQR and GCCQR using only single attribute groups as features in the optimization problem, excluding 2-way combinations. of attributes and the values they take on. In total, we have $|\mathcal{G}|=77$ subgroups. [^6] To prevent extremely uncommon groups that may exist in the Wikidata database from biasing our results, we exclude groups of size $<5\%$ of the total test set size.
4.3 Generating confidence scores
As mentioned previously, the algorithms described in Section 3 require a base scoring function. For our experiments, we use a frequency-based scoring function derived from concept of self-consistency (Wang et al., 2022). Specifically, to score each claim found in a generated biography, we prompt the LLM to output a biography $M$ additional times. We use the proportion of times the claim is contained in the additional reference generations as the uncertainty score. To automate this process, we use a procedure similar to frequency scoring algorithm proposed by Wang et al. (2024) in which (1) a set of $K$ most relevant claims from a reference generation is retrieved using a vanilla BM25 algorithm (to reduce computational costs). Then (2) an LLM is tasked to evaluate whether the target claim is supported by the set of $K$ reference claims. In our work, we replace LLM prompting in step (2) with AlignScore-Large (Zha et al., 2023), which runs significantly faster and is reported by Zha et al. (2023) to compare favorably to LLM-based alignment methods.
5 Results
To assess the efficacy of these methods for quantifying the uncertainty of LLM factuality, we present empirical results for the task of biography generation, using outputs from Llama 2 7B Chat (Touvron et al., 2023) and Mistral 7B Instruct v0.2 (Jiang et al., 2023). We randomly split the entities into 80-20 calibration-test splits. The results we present are averaged over 10 distinct randomly generated splits. Table 5 reports the number of entities for which each model generates biographies and the number of claims contained in each generated biography.
Calibration.
To evaluate fact-level uncertainty, we consider both the ASCE and the Brier score, which is the mean squred error between the uncertainty score function $f(X)$ and the true label $Y$. While not a direct measure of (multi)calibration like the ASCE, the Brier score is still useful in certain settings for quantifying the efficacy of the algorithms we consider. As discussed previously, we generate biographies for entities from Bio-NQ and apply the techniques discussed in Section 3.
In Table 1, we report the ASCE, max gASCE, and mean gASCE, comparing each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). The multicalibration variants of both the patching and linear regression techniques significantly outperform HB and PS in terms of max and mean gASCE. More surprising, however, is that even when considering just the ASCE across the entire dataset, incorporating group features improves performance as well, especially when comparing IGHB to HB. These results suggest that even if one does not specifically require parity for specific subgroups, collecting additional group features and applying multicalibration (as opposed to vanilla calibration) can still be extremely beneficial for generating better-calibrated uncertainty scores. In Table 2, we report Brier score, also both marginally and across groups. Again, we observe that IGHB and GCULR outperform HB and PS respectively across all metrics.
Conformal prediction.
For the problem of uncertainty at the biography level, we apply the vanilla conformal prediction methods SC and CQR and their multivalid counterparts, MVSC and GCCQR. We choose target coverages of between $0.5$ to $0.9$. [^7] Although we evaluate on a wide set of target coverages $1-\alpha$, conformal prediction makes more sense only for higher target coverages (e.g., $0.8$ or higher), since lower coverage guarantees can often be too weak to be useful in practice. We initially observe that for both Llama 2 7B Chat and Mistral 7B Instruct, all methods are able to achieve close to perfect coverage on Bio-NQ (Figure 2, left panel). However, when evaluating coverage across individual subgroups, we find that all methods have some level of error. In Figure 1, we compare the mean absolute coverage across all subgroups for each target coverage. Like for calibrating individual facts, we again find that multivalid variants of both the patching (MVSC) and linear regression (GCCQR) techniques outperform standard conformal techniques (SC, CQR).
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 1: For each target coverage $1-\alpha$, we run conformal prediction methods (SC, CQR) and their multigroup counterparts (MVSC, GCCQR). On the top row of plots (a), we evaluate on biographies generated by Llama 2 7B Chat, and on the bottom row (b), we evaluate on biographies generated by Mistral 7B Instruct. We calculate the average coverage error across all subgroups and plot them side by side for each pair of methods.
6 Conclusion
In this paper, we conduct an extensive study on uncertainty quantification in long-form LLM generation. We focus on two forms of uncertainty quantification for factuality in long-form generation—claim-level (calibration) and biography-level (conformal prediction)—and present a variety of methods for these settings. Introducing two categories of algorithms (iterative patching and linear regression), we demonstrate that by accounting for additional groups, multicalibration and multivalid conformal prediction methods outperform their marginal-guarantee counterparts across group-wise performance and, in the case of calibration, overall marginal performance. We consider these empirical results to establish a benchmark for this setting and hope our findings will motivate future work in quantifying uncertainty for LLMs.
7 Limitations and potential risks
Past works on uncertainty quantification for NLP have made great strides in providing confidence scores for text generation, and our work furthers this line of research. However, we note that there exists certain limitations in our empirical findings. First, while Min et al. (2023a) demonstrate the effectiveness of using an LLM to both decompose generations into atomic facts and assess the factuality of these claims, this automated process serves only as a proxy for the gold standard of human annotation. Moreover, like in Min et al. (2023a), our assessment of factuality is limited to whether claims are supported by Wikipedia, which again serves only as a proxy for factuality.
In our experiments, we chose groups based on attributes that are general and easily obtainable for any public figure (i.e., Wikidata properties). This set of groups is non-exhaustive, and there may be many other sets of groups one can consider that may be relevant for various problem domains.
Finally, while we follow prior work and focus on the problem of generating biographies, our uncertainty techniques extend broadly to any long-form generation task where information contained in the output can be broken down into independent claims. Furthermore, the automated evaluation pipeline proposed by Min et al. (2023a) is still valid, as long as the entities one prompts the LLM for a description of can be found on Wikipedia (i.e., animals, places, historical events, etc.). Consequently, in future work, we hope to validate our findings for other long-form generation tasks.
Regarding potential risks, as mentioned previously, large language models can have a tendency to hallucinate. Moreover, deep learning models in general may exhibit or perpetuate bias. In an attempt to better quantify uncertainty of LLMs, our work aims to help alleviate the potential risks of models producing false information. Moreover, our emphasis on multi-group guarantees aligns with certain definitions of fairness. However, our work may not entirely mitigate such issues, and other potential malicious or unintended harmful effects could still persist in the outputs of the calibration and conformal methods we evaluate.
Appendix A Details on biography generation and factuality evaluation
While the ground truth score must be human-annotated, Min et al. (2023a) show that FActSCORE can be approximated by an automated process that leverages an LLM (i.e., ChatGPT and LLaMa-7B) and natural language retrieval. Following Min et al. (2023a), we also use an LLM to automate the annotation process. For some input person, [ENTITY], we prompt a large language model with the following:
[INST] Question: Tell me a bio of [ENTITY]. [/INST]
We then decompose each long-form generation into a set of atomic facts, which are then checked against some set of Wikipedia articles about the [ENTITY] to evaluate overall performance of language model in terms of factuality. Min et al. (2023a) demonstrate that while the evaluation process should ideally be conducted by human annotators, using large language models (i.e., ChatGPT and LLama 1 7B) to both decompose long-form generations and check against Wikipedia articles serves as a very good proxy for human annotation.
For checking each atomic fact against Wikpedia, we directly use the code released by Min et al. (2023a), which first conducts passage retrieval via Generalizable T5-based Retrievers (Liu et al., 2023) to find relevant articles from a dump of Wikipedia (dated 2023-04-01) and then prompts an LLM (i.e., ChatGPT or Llama 1 7B) to predict whether each fact is supported by the retrieved passages. For our evaluation, we again use Llama 2 7B Chat. Finally, these predictions are ensembled with predictions using likelihood estimates derived from a nonparametric masked language model (Min et al., 2023b).
Appendix B Additional results
B.1 Additional results on Bio-NQ
In Figure 2, we provide additional information about the prediction sets outputted by our various conformal methods on Bio-NQ. On the left panel, we plot the empirical coverage achieved against the target coverage. Figure 2 demonstrates that all methods achieve the target (marginal) coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level, while on the right panel, we plot the number of facts per biography retained. Generally speaking, all methods retain about the same number of facts per biography. We also observe that to achieve the same target coveragre, SC and MVSC generally retain fewer biographies (i.e., more abstentions) when compared to CQR and GCCQR. However, when comparing each conformal method (SC and CQR) to their multivalid counterparts (MVSC and GCCQR), we again observe that there are very little differences between them.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 2: We report additional metrics for conformal predictions techniques when evaluated on biographies generated for Bio-NQ: On the top row, we present these metrics for outputs from LLama 2 7B Chat, and on the bottom, Mistral 7B Instruct. On the left panel, we plot the empirical coverage achieved against the target coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level. Finally, on the right panel, we plot the number of facts per biography retained, again against the target coverage level.
B.2 Evaluating on entities used in Min et. al (2023a)
In addition to evaluating on our dataset, Bio-NQ, we construct an additional dataset using the $683$ entities used by Min et al. (2023a) for their empirical evaluation. We denote this dataset as Bio-FActScore.
Calibration.
In Table 3, we observe similar results to that on Bio-NQ—namely, multicalibrated counterparts (IGHB and GCULR) perform better than their base variant (HB and PS). However, we note that for Mistral 7B Instruct, PS performs the best when looking at marginal ASCE. We hypothesize that the smaller gap in ASCE between PS and GCULR may be due to the smaller training size of Bio-FActScore (25,283 claims), which is roughly 10x smaller than that of Bio-NQ (297,714 claims). Lastly, with respect to Brier score, multicalibration still dominates across all metrics (Table 4).
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.26830 0.00951 0.00229 0.00164 0.00125*
group max 0.48594 0.07208 0.04088 0.05017 0.03519*
group mean 0.29983 0.02848 0.01108 0.01659 0.00858*
group max 0.54701 0.08436 0.04585* 0.07043 0.04931
group mean 0.29435 0.03226 0.01143 0.01848 0.00911*
Table 3: We generate biographies for entities from Bio-FActScore. We evaluate using ASCE, max gASCE, and average gASCE ($\downarrow$ better), comparing each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better performing method for each pairing. * denotes the best performing method across all methods evaluated.
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.475 0.169 0.148 0.152 0.143*
group max 0.535 0.323 0.247 0.285 0.235*
group mean 0.479 0.169 0.148 0.152 0.142*
Mistral 7B Instruct marginal 0.471 0.186 0.159 0.164 0.152*
group max 0.554 0.333 0.250 0.285 0.235*
group mean 0.477 0.186 0.158 0.164 0.152*
Table 4: We generate biographies for entities from Bio-FActScore. We evaluate using Brier score ($\downarrow$ better) marginally across the entire dataset, as well as within each subgroup (in terms of max and mean overall all groups). We compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better performing method for each pairing. * denotes the best performing method across all methods evaluated.
Conformal Prediction.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 3: For each target coverage $1-\alpha$, we run conformal prediction methods (SC, CQR) and their multigroup counterparts (MVSC, GCCQR). On the top row of plots (a), we run the methods on biographies generated by Llama 2 7B Chat, and on the bottom row (b), we run on biographies generated by Mistral 7B Instruct. We calculate the average coverage error across all subgroups and plot them side by side for each pair of methods.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 4: We report additional metrics for conformal predictions techniques when evaluated on biographies generated for Bio-FActScore: On the top row, we present these metrics for outputs from LLama 2 7B Chat, and on the bottom, Mistral 7B Instruct. On the left panel, we plot the empirical coverage achieved against the target coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level. Finally, on the right panel, we plot the number of facts per biography retained, again against the target coverage level.
Appendix C Additional experimental details
Code for reproducing all results will accompany the final version of this work.
Datasets.
dataset model # entities total # claims avg. # claims per bio.
Bio-FAcTScore Llama 2 7B Chat 683 17,605 25.78
Mistral 7B Instruct 683 25,283 37.02
Bio-NQ Llama 2 7B Chat 8,541 206,620 24.19
Mistral 7B Instruct 8,541 297,714 34.86
Table 5: We provide statistics about our two datasets and how many claims are generated by each LLM.
@@ Appendix C Additional experimental details
For AlignScore, we set $M=4$ and $K=5$. We found that AlignScore generally returns values close to 00 or $1$, giving us uncertainty scores around the $5$ values $\{0,\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}$. As a result, we evaluate all methods using $p=5$ level sets. We note that instead of rounding to $\{0,\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}$, we use a $k$-means clustering algorithm, where $k=5$, on the calibration set to determine what values to round to for IGHB (after each iteration), PS, and GCULR. Note that this method of rounding is not applicable to HB.
@@
Wikidata and Wikipedia are licensed under the Creative Commons CC0 License. Llama 2 7B is licensed under Meta’s Llama 2 license. Mistral 7B Chat and Hugging Face’s transformers library are licensed under Apache 2.0 license. We also make use of code released by Min et al. (2023a) under the MIT license.

### v2

@@ B.3 Miscellaneous details
  • nationality: Following Min et al. [2023a], who use nationality derived from Wikidata to sample their dataset of human entities, we take the property country of citizenship (or place of birth when not available) and categorize the corresponding value into the following categories defined by Min et al. [2023a]: Asia/Pacific, Europe/Middle East, North America, Latin/South America/Africa.
@@
In total, we have $|\mathcal{G}|=77$ subgroups. To prevent extremely uncommon groups that may exist in the Wikidata database from biasing our results, we exclude groups of size $<5\%$ of the total test set size. Note that while we create groups using $1$ and $2$-way combinations for evaluation, we train the quantile regression models in CQR and GCCQR using only single attribute groups as features in order to reduce computation.
@@
For AlignScore, we set $M=4$ and $K=5$. We found that AlignScore generally returns values close to 00 or $1$, giving us self-consistency uncertainty scores around the $5$ values $\{0,\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}$. As a result, we evaluate all methods using $p=5$ level sets.
@@
Wikidata and Wikipedia are licensed under the Creative Commons CC0 License. Llama 2 7B is licensed under Meta’s Llama 2 license. Mistral 7B Chat and Hugging Face’s transformers library are licensed under Apache 2.0 license. We also make use of code released by Min et al. [2023a] under the MIT license.
Appendix C Evaluating on entities used in Min et. al (2023a)
In addition to evaluating on our dataset, Bio-NQ, we construct an additional dataset using the $683$ entities used by Min et al. [2023a] for their empirical evaluation. We denote this dataset as Bio-FActScore and evaluate all methods using self-consistency as the base scoring function.
Calibration.
In Table 16, we observe similar results to that on Bio-NQ—namely, multicalibrated counterparts (IGHB and GCULR) perform better than their base counterpart (HB and PS). However, we note that for Mistral 7B Instruct, PS performs the best when looking at marginal ASCE. We hypothesize that the smaller gap in ASCE between PS and GCULR may be due to the smaller training size of Bio-FActScore (25,283 claims), which is roughly 10x smaller than that of Bio-NQ (297,714 claims). Lastly, with respect to Brier score, multicalibration still dominates across all metrics (Table 17).
Model Metric Uncalibrated HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.26830 0.00951 0.00229 0.00164 0.00125*
group max 0.48594 0.07208 0.04088 0.05017 0.03519*
group mean 0.29983 0.02848 0.01108 0.01659 0.00858*
group max. 0.54701 0.08436 0.04585* 0.07043 0.04931
group mean 0.29435 0.03226 0.01143 0.01848 0.00911*
Table 16: We generate biographies for entities from Bio-FActScore and compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR) on ASCE, max gASCE, and average gASCE ($\downarrow$ better). We bold the better-performing method for each pairing. * denotes the best-performing method across all methods evaluated. All methods use self-consistency as their base scoring function.
Model Metric Uncalibrated HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.475 0.169 0.148 0.152 0.143*
group max 0.535 0.323 0.247 0.285 0.235*
group mean 0.479 0.169 0.148 0.152 0.142*
Mistral 7B Instruct marginal 0.471 0.186 0.159 0.164 0.152*
group max 0.554 0.333 0.250 0.285 0.235*
group mean 0.477 0.186 0.158 0.164 0.152*
Table 17: We generate biographies for entities from Bio-FActScore and compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR) on Brier score ($\downarrow$ better) marginally across the entire dataset, as well as within each subgroup (in terms of max and mean over all groups). We bold the better-performing method for each pairing. * denotes the best-performing method across all methods evaluated. All methods use self-consistency as their base scoring function.
Conformal Prediction.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 3: For each target coverage, we run conformal methods (SC, CQR) and their multigroup counterparts (MVSC, GCCQR) on Bio-FactScore. We evaluate on generations by (a) Llama 2 7B Chat and (b) Mistral 7B Instruct. We calculate the average coverage error across all groups and plot them side by side for each pairing. All methods use self-consistency as their base scoring function.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 4: We report additional metrics for conformal predictions techniques using self-consistency scores when evaluated on biographies generated for Bio-FActScore: On the top row, we present these metrics for outputs from LLama 2 7B Chat, and on the bottom, Mistral 7B Instruct. On the left panel, we plot the empirical coverage achieved against the target coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level. Finally, on the right panel, we plot the number of facts per biography retained, again against the target coverage level.